### v1

@@ Abstract
Human body restoration, as a specific application of image restoration, is widely applied in practice and plays a vital role across diverse fields. However, thorough research remains difficult, particularly due to the lack of benchmark datasets. In this study, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline. This pipeline leverages existing object detection datasets and other unlabeled images to automatically crop and filter high-quality human images. Using this pipeline, we constructed a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset, which includes training, validation, and test sets. The dataset significantly surpasses other human-related datasets in both quality and content richness. Finally, we propose OSDHuman, a novel one-step diffusion model for human body restoration. Specifically, we propose a high-fidelity image embedder (HFIE) as the prompt generator to better guide the model with low-quality human image information, effectively avoiding misleading prompts. Experimental results show that OSDHuman outperforms existing methods in both visual quality and quantitative metrics. The dataset and code will at https://github.com/gobunu/OSDHuman.
@@ 1 Introduction
Human body restoration (HBR) aims to recover high-quality (HQ) images from low-quality (LQ) inputs featuring human figures. Unlike nature scene pictures, human figures naturally attract viewers’ attention in images. However, real-world images often suffer from degradation during capture and transmission, such as blur, noise, resolution reduction, and JPEG artifacts. These distortions severely impact the recognition of human activities and the extraction of information from the image. Furthermore, degraded images make other human-related downstream tasks more challenging, such as human-object interaction detection Wang et al. (2024d), human pose estimation Samkari et al. (2023), and 3D reconstruction Wang et al. (2021a).
Despite its practical significance, progress in HBR remains constrained, primarily due to the absence of task-specific benchmark datasets. In natural scenarios, humans exhibit a wide range of activities and complex interactions with their surroundings. Therefore, the benchmark dataset for HBR must be large, cover complex scenarios, and include natural activities. Datasets in the fashion domain, such as DeepFashion Liu et al. (2016) and iDesigner Dufour et al. (2022), focus on runway or studio scenarios. As a result, these datasets contain only limited types of human actions, making them unsuitable for HBR. Moreover, human images often involve multiple individuals interacting with each other, adding further complexity to the restoration task. Such complexity makes datasets focused on single-person image generation, such as SHHQ Fu et al. (2022) and CosmicMan-HQ Li et al. (2024b), less suitable for HBR, as they mainly focus on single-person images.
Refer to caption Refer to caption Refer to caption Refer to caption LQ (512$\times$512) OSEDiff SinSR ResShift Refer to caption Refer to caption Refer to caption Refer to caption OSDHuman OSEDiff* Sinsr* ResShift* (ours) Wu et al. (2024a) Wang et al. (2024e) Yue et al. (2023)
@@
Additionally, an HBR model can achieve optimal performance only when trained on sufficiently high-quality datasets. Degraded datasets could cause bias in the model’s weights, as it is difficult to distinguish between degradation and features in LQ images. Some existing image restoration datasets, such as LSDIR Li et al. (2023) and DIV2K Agustsson & Timofte (2017), are of high quality and cover complex real-world scenarios. However, the proportion of human images in these datasets is small, and they are not specifically tailored for human images. Other human-related high-level datasets, such as those for object detection Kuznetsova et al. (2020); Shao et al. (2019) and keypoint detection Lin et al. (2014), have a wider range of human activities and sophisticated surrounding objects due to the diversity of image sources. However, as illustrated in Fig. 1, these datasets lack dedicated quality filtering, containing substantial LQ samples.
@@
However, even with a high-quality benchmark, achieving excellent HBR performance still requires a well-designed model architecture. In recent studies, image restoration models with latent diffusion model (LDM) Rombach et al. (2022) architecture achieve promising results due to powerful generative capabilities. These models combine generation and restoration to reconstruct lost parts of LQ images using the features provided. They primarily use two main latent space mapping methods: variational autoencoder (VAE) Kingma & Welling (2014) and vector quantized VAE (VQVAE) Oord et al. (2017). As shown in Fig. 2, models with VQVAE, such as ResShift Yue et al. (2023) and SinSR Wang et al. (2024e), often produce distorted structures in detail. Even with retraining, these issues are difficult to resolve due to the VQVAE codebook’s inability to capture the fine details of the human body. On the other hand, models with VAE, like OSEDiff Wu et al. (2024a), have better generalization and detail generation capabilities, but they are still not specifically optimized for HBR.
@@
While multi-step diffusion models have strong restoration abilities for LQ images, they often require substantial computational resources, which limits their applicability. To reduce resource consumption, one-step diffusion (OSD) models are proposed and achieve good results. By leveraging large-scale pretrained text-to-image (T2I) models Saharia et al. (2022); Rombach et al. (2022) as foundation models, OSD models combine generation power with fast inference. Therefore, OSD models are highly competitive in HBR. However, this also requires the model to incorporate an appropriate prompt extractor that can derive a high-fidelity prompt from complex human images. Otherwise, the resulting prompt could mislead the restoration of the model.
@@
To address the limitations, we propose OSDHuman, a novel OSD model for HBR. Firstly, to overcome the lack of benchmark datasets in HBR, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline. This pipeline preprocesses both labeled and unlabeled datasets to isolate images containing humans. Then it refines human bounding boxes and crops them accordingly. Using no-reference image quality assessment (IQA) metrics, we ultimately produce a dataset. Secondly, leveraging HQ-ACF, we develop a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset, which comprises 109,052 HQ 512$\times$512 human images for training. This pipeline also provides images for validation and testing. The PERSONA dataset includes both individual-environment interactions and multi-person interactions, averaging 3.4154 individuals per image. Thirdly, to provide prompts suitable for HBR, we propose a high-fidelity image embedder (HFIE). HFIE uses an image encoder from RAM Zhang et al. (2023) and a multi-head attention layer with a learnable embedding as the query. This design avoids distortions introduced by tags when summarizing images, thereby preventing misleading prompts that could impair model restoration. In addition, we employ a variational score distillation (VSD) regularizer to guide the model’s generative distribution toward natural image distributions.
@@
Figure 3: High-quality dataset automated cropping and filtering pipeline. The pipeline consists of four stages. First, multiple datasets are collected, comprising millions of images. Images without labels are processed using YOLO11 for human detection. Then, a Laplacian operator is applied to compute image Laplacian variance, filtering out images below a threshold. Next, human boxes are adjusted to the square shape, and overly small or densely packed boxes are removed. Finally, cropped human images are evaluated using Image Quality Assessment (IQA) metrics. Images ranking in the top third by normalized metrics and exceeding the metric threshold are selected. These 109,052 images constitute the person-based restoration with sophisticated objects and natural activities (PERSONA) dataset.
@@
  • Our OSDHuman achieves strong SOTA human body restoration performance, excelling in visual quality and metrics while maintaining lower computational costs.
@@ 2.1 Human Body restoration
Human body restoration could benefit both the fashion industry for display and the photography and camera producers. The purpose is evident, focusing on the human body and making it look better. Compared to general image super-resolution (SR) for arbitrary objects, human body image SR is more constrained, allowing for the use of a variety of prior knowledge. Firstly, current research on image segmentation has made significant progress in generating segmentation masks for different body parts, such as hair and arms, which provide a clear description of the human body shape. This knowledge is beneficial for handling the boundaries in the human body SR tasks. Since the human body composition is relatively fixed and limited, it is easier to estimate compared to the random and arbitrary objects in natural images.
Lots of research has been done recently. Liu et al. (2021a) represent the texture details of the human body using subbands of the non-subsampled shearlet transform, while Wang et al. (2024c) employ a pyramid residual network to estimate texture and shape priors, improving the image quality of the human body. Zhang et al. (2024), as the first to apply diffusion models for enhancing body-region images, use pose-attention guidance, text guidance, and a body-centered diffusion sampler to integrate various semantic information for high-quality body-region enhancement.
@@ 2.2 Diffusion Models
Since the diffusion model was released and popular, many efforts have been made. Two classical applications are image restoration and text-to-image (T2I). Image restoration, as the first and most natural application, has been developed a lot Saharia et al. (2023); Whang et al. (2022); Avrahami et al. (2022); Chen et al. (2023); Xia et al. (2023). With the development of conditional diffusion models Rombach et al. (2022), numerous companies have invested heavily in training more powerful T2I models. Recently, many efforts have been made to integrate these two typical applications. The rapidly developing Stable Diffusion Rombach et al. (2022), DALLE Ramesh et al. (2021), and PixArt Chen et al. (2024) have continually pushed the boundaries of realism and diversity in T2I generation. Many image restoration methods have also leveraged pretrained models to achieve more natural image recovery Wu et al. (2024b); Yang et al. (2023); Lin et al. (2024); Wu et al. (2024a); Wang et al. (2024a).
Stricted on the multi-timesteps in the diffusion inference procedure, the diffusion models with 50 or more steps Wang et al. (2024b); Lin et al. (2024); Wu et al. (2024b); Yang et al. (2023) cannot actually be used in practice. Many efforts have been made to faster diffusion models, such as cutting, quantizing, and compressing. Moreover, eliminating the number of inference timesteps is a convincing way, especially applied in image restoration. SinSR Wang et al. (2024e) pioneers one-step inference for diffusion-based SR by distilling deterministic generation functions into a student network, coupled with a consistency-preserving loss and efficient training pair generation strategy. OSEDiff Wu et al. (2024a) adapts pretrained SD models for Real-ISR through LoRA-finetuned U-Net and variational score distillation, enabling direct low-quality image reconstruction in one step without noise injection. Those methods achieve a fascinating performance in natural image restoration.
@@ 3.1 High Quality Human Dataset Pipeline
For image restoration tasks, large-scale high-quality datasets are required to simulate various real-world scenarios and objects. There are already many high-quality image restoration datasets, such as FFHQ Karras et al. (2019) and LSDIR Li et al. (2023). However, these datasets are not ideally suited for the human body restoration (HBR) task due to their lack of focus on human-specific features. Moreover, the dataset should enable models to adapt to real-world environments’ complexities. It must encompass most scenarios including interactions among people and between people and their surroundings. Therefore, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline for HBR datasets, as well as a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset.
@@
Automated Cropping and Filtering Pipeline. As illustrated in Fig. 3, we first collect a series of commonly used and publicly available large-scale object detection datasets, including COCO Lin et al. (2014), OID Kuznetsova et al. (2020); Krasin et al. (2017), Object365 Shao et al. (2019) and CrowdHuman Shao et al. (2018), comprising approximately 4 million images. We then filter the images by labels, selecting those containing “human” or synonymous labels, such as “Human Body” in OID and “person” in Object365. To further refine the selection, we conduct human detection on the image restoration dataset LSDIR Li et al. (2023), using the YOLO11 model Jocher & Qiu (2024) for processing. This operation resulted in bounding boxes similar to those in object detection datasets.
@@
Next, we apply the Laplacian operation to these datasets and compute the variance of the results. Images with a variance below the threshold are discarded, as these correspond to images with a high degree of blurriness. Before cropping, we also check the size of the human bounding boxes. Images with low-resolution human bodies are rejected. After these steps, we use the bounding boxes to crop the images. When cropping, the side length of the bounding box’s longer edge is used as the side length of the cropping box, ensuring a square crop. In cases where an image contains multiple overlapping boxes, non-maximum suppression (NMS) is applied, prioritizing the box closest to the image center. The cropped images are then resized to 512$\times$512.
@@
Finally, we obtain approximately 440,000 cropped images, on which we measure no-reference Image Quality Assessment (IQA) metrics. The metrics used are common in image restoration, including CLIPIQA Wang et al. (2023a), MANIQA Yang et al. (2022), MUSIQ Ke et al. (2021), and NIQE Zhang et al. (2015). To balance the evaluation performance of each metric, we normalize the obtained IQA metrics using the following standardization formula:
Normalized $\text{Normalized Metrics}=\frac{1}{N}\sum_{i=1}^{N}\frac{M_{i}-\mu_{i}}{\sigma_{i}},$ (1)
@@
where $\mu_{i}$ is the mean and $\sigma_{i}$ is the standard deviation of the metrics ($M_{i}$). Since NIQE is better when its score is smaller, we first apply a negative transformation to its values. The normalized scores are then accumulated per image and sorted. The final PERSONA dataset version consists of 109,052 images with normalized metrics in the top third, and each IQA metric must exceed a predefined threshold.

### v2

@@ Abstract
Human body restoration, as a specific application of image restoration, is widely applied in practice and plays a vital role across diverse fields. However, thorough research remains difficult, particularly due to the lack of benchmark datasets. In this study, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline. This pipeline leverages existing object detection datasets and other unlabeled images to automatically crop and filter high-quality human images. Using this pipeline, we constructed a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset, which includes training, validation, and test sets. The dataset significantly surpasses other human-related datasets in both quality and content richness. Finally, we propose OSDHuman, a novel one-step diffusion model for human body restoration. Specifically, we propose a high-fidelity image embedder (HFIE) as the prompt generator to better guide the model with low-quality human image information, effectively avoiding misleading prompts. Experimental results show that OSDHuman outperforms existing methods in both visual quality and quantitative metrics. The dataset and code are available at: https://github.com/gobunu/OSDHuman.
Machine Learning, ICML
@@ 1 Introduction
Human body restoration (HBR) aims to recover high-quality (HQ) images from low-quality (LQ) inputs featuring human figures. Unlike nature scene pictures, human figures naturally attract viewers’ attention in images. However, real-world images often suffer from degradation during capture and transmission, such as blur, noise, resolution reduction, and JPEG artifacts. These distortions severely impact the recognition of human activities and the extraction of information from the image. Furthermore, degraded images make other human-related downstream tasks more challenging, such as human-object interaction detection (Wang et al., 2024e; Liu et al., 2025), human pose estimation (Samkari et al., 2023; Atmosukarto et al., 2024), and 3D reconstruction (Wang et al., 2021a; Sun et al., 2024).
Despite its practical significance, progress in HBR remains constrained, primarily due to the absence of task-specific benchmark datasets. In natural scenarios, humans exhibit a wide range of activities and complex interactions with their surroundings. Therefore, the benchmark dataset for HBR must be large, cover complex scenarios, and include natural activities. Datasets in the fashion domain, such as DeepFashion (Liu et al., 2016) and iDesigner (Dufour et al., 2022), focus on runway or studio scenarios. As a result, these datasets contain only limited types of human actions, making them unsuitable for HBR. Moreover, human images often involve multiple individuals interacting with each other, adding further complexity to the restoration task. Such complexity makes datasets focused on single-person image generation, such as SHHQ (Fu et al., 2022) and CosmicMan-HQ (Li et al., 2024b), less suitable for HBR, as they mainly focus on single-person images.
Refer to caption Refer to caption Refer to caption Refer to caption LQ ($\times512$) OSEDiff SinSR ResShift Refer to caption Refer to caption Refer to caption Refer to caption OSDHuman OSEDiff* SinSR* ResShift* (ours) (Wu et al., 2024a) (Wang et al., 2024d) (Yue et al., 2023)
@@
Additionally, an HBR model can achieve optimal performance only when trained on sufficiently high-quality datasets. Degraded datasets could cause bias in the model’s weights, as it is difficult to distinguish between degradation and features in LQ images. Some existing image restoration datasets, such as LSDIR (Li et al., 2023) and DIV2K (Agustsson and Timofte, 2017), are of high quality and cover complex real-world scenarios. However, the proportion of human images in these datasets is small, and they are not specifically tailored for human images. Other human-related high-level datasets, such as those for object detection (Kuznetsova et al., 2020; Shao et al., 2019) and keypoint detection (Lin et al., 2014), have a wider range of human activities and sophisticated surrounding objects due to the diversity of image sources. However, as illustrated in Fig. 1, these datasets lack dedicated quality filtering, containing substantial LQ samples.
@@
However, even with a high-quality benchmark, achieving excellent HBR performance still requires a well-designed model architecture. In recent studies, image restoration models with latent diffusion model (LDM) (Rombach et al., 2022) architecture achieve promising results due to powerful generative capabilities. These models combine generation and restoration to reconstruct lost parts of LQ images using the features provided. They primarily use two main latent space mapping methods: variational autoencoder (VAE) (Kingma and Welling, 2014) and vector quantized VAE (VQVAE) (Oord et al., 2017). As shown in Fig. 2, models with VQVAE, such as ResShift (Yue et al., 2023) and SinSR (Wang et al., 2024d), often produce distorted structures in detail. Even with retraining, these issues are difficult to resolve due to the VQVAE codebook’s inability to capture the fine details of the human body. On the other hand, models with VAE, like OSEDiff (Wu et al., 2024a), have better generalization and detail generation capabilities, but they are still not specifically optimized for HBR.
@@
While multi-step diffusion models have strong restoration abilities for LQ images, they often require substantial computational resources, which limits their applicability. To reduce resource consumption, one-step diffusion (OSD) models are proposed and achieve good results. By leveraging large-scale pretrained text-to-image (T2I) models (Saharia et al., 2022; Rombach et al., 2022) as foundation models, OSD models combine generation power with fast inference. Therefore, OSD models are highly competitive in HBR. However, this also requires the model to incorporate an appropriate prompt extractor that can derive a high-fidelity prompt from complex human images. Otherwise, the resulting prompt could mislead the restoration of the model.
@@
To address the limitations, we propose OSDHuman, a novel OSD model for HBR. Firstly, to overcome the lack of benchmark datasets in HBR, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline. This pipeline preprocesses both labeled and unlabeled datasets to isolate images containing humans. Then it refines human bounding boxes and crops them accordingly. Using no-reference image quality assessment (IQA) metrics, we ultimately produce a dataset. Secondly, leveraging HQ-ACF, we develop a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset, which comprises 109,053 HQ $\times512$ human images for training. This pipeline also provides images for validation and testing. The PERSONA dataset includes both individual-environment interactions and multi-person interactions, averaging 3.4154 individuals per image. Thirdly, to provide prompts suitable for HBR, we propose a high-fidelity image embedder (HFIE). HFIE uses an image encoder from RAM (Zhang et al., 2023) and a multi-head attention layer with a learnable embedding as the query. This design avoids distortions introduced by tags when summarizing images, thereby preventing misleading prompts that could impair model restoration. In addition, we employ a variational score distillation (VSD) regularizer to guide the model’s generative distribution toward natural image distributions.
@@
Figure 3: High-quality dataset automated cropping and filtering pipeline. The pipeline consists of four stages. First, multiple datasets are collected, comprising millions of images. Images without labels are processed using YOLO11 for human detection. Then, a Laplacian operator is applied to compute image Laplacian variance, filtering out images below a threshold. Next, human boxes are adjusted to the square shape, and overly small or densely packed boxes are removed. Finally, cropped human images are evaluated using Image Quality Assessment (IQA) metrics. Images ranking in the top third by normalized metrics and exceeding the metric threshold are selected. These 109,053 images constitute the person-based restoration with sophisticated objects and natural activities (PERSONA) dataset.
@@
  • Our OSDHuman achieves state-of-the-art human body restoration performance, excelling in visual quality and metrics while maintaining lower computational costs.
@@ 2.1 Human Body Restoration
Human body restoration (HBR) could benefit both the fashion industry for display and the photography and camera producers. The purpose is evident, focusing on the human body and making it look better. Compared to general image restoration for arbitrary objects, HBR is more constrained, allowing for the use of a variety of prior knowledge. Firstly, current research on image segmentation has made significant progress in generating segmentation masks for different body parts, such as hair and arms, which provide a clear description of the human body shape. This knowledge is beneficial for handling the boundaries in the HBR tasks. Since the human body composition is relatively fixed and limited, it is easier to estimate compared to the random and arbitrary objects in natural images.
Lots of research has been done recently. A previous work (Liu et al., 2021a) captures body texture using subbands of the non-subsampled shearlet transform, while PRCN (Wang et al., 2024c) employs a pyramid residual network to estimate texture and shape priors, enhancing body images. DiffBody (Zhang et al., 2024), as the first to apply diffusion models, uses pose-attention, text guidance, and a body-centered sampler to integrate semantic information for body-region enhancement.
@@ 2.2 Diffusion Models
Since the diffusion model was released and popular, many efforts have been made. Two classical applications are image restoration and text-to-image (T2I). Image restoration, as the first and most natural application, has been developed a lot (Saharia et al., 2023; Whang et al., 2022; Avrahami et al., 2022; Chen et al., 2023; Xia et al., 2023). With the development of conditional diffusion models (Rombach et al., 2022), numerous companies have invested heavily in training more powerful T2I models. Recently, many efforts have been made to integrate these two typical applications. The rapidly developing Stable Diffusion (Rombach et al., 2022), DALLE (Ramesh et al., 2021), and PixArt (Chen et al., 2024) have continually pushed the boundaries of realism and diversity in T2I generation. Many image restoration methods have also leveraged pretrained models to achieve more natural image recovery (Wu et al., 2024b; Yang et al., 2024; Lin et al., 2024; Wu et al., 2024a; Wang et al., 2024b).
Stricted on the multi-step in the diffusion inference procedure, the diffusion models with 50 or more steps (Wang et al., 2024a; Lin et al., 2024; Wu et al., 2024b; Yang et al., 2024) cannot actually be used in practice. Many efforts have been made to faster diffusion models, such as cutting, quantizing, and compressing. Moreover, eliminating the number of inference timesteps is a convincing way, especially applied in image restoration. SinSR (Wang et al., 2024d) pioneers one-step inference for diffusion-based super-resolution (SR) by distilling deterministic generation functions into a student network, coupled with a consistency-preserving loss and efficient training pair generation strategy. OSEDiff (Wu et al., 2024a) adapts pretrained SD models for SR through LoRA-finetuned U-Net and variational score distillation, enabling direct low-quality image reconstruction in one step without noise injection. Those methods achieve a fascinating performance in natural image restoration.
@@ 3.1 High Quality Human Dataset Pipeline
For image restoration tasks, large-scale high-quality datasets are required to simulate various real-world scenarios and objects. There are already many high-quality image restoration datasets, such as FFHQ (Karras et al., 2019) and LSDIR (Li et al., 2023). However, these datasets are not ideally suited for the human body restoration (HBR) task due to their lack of focus on human-specific features. Moreover, the dataset should enable models to adapt to real-world environments’ complexities. It must encompass most scenarios including interactions among people and between people and their surroundings. Therefore, we propose a high-quality dataset automated cropping and filtering (HQ-ACF) pipeline for HBR datasets, as well as a person-based restoration with sophisticated objects and natural activities (PERSONA) dataset.
@@
Automated Cropping and Filtering Pipeline. As illustrated in Fig. 3, we first collect a series of commonly used and publicly available large-scale object detection datasets, including COCO (Lin et al., 2014), OID (Kuznetsova et al., 2020; Krasin et al., 2017), Object365 (Shao et al., 2019) and CrowdHuman (Shao et al., 2018), comprising approximately 4 million images. We then filter the images by labels, selecting those containing “human” or synonymous labels, such as “Human Body” in OID and “person” in Object365. To further refine the selection, we conduct human detection on the image restoration dataset LSDIR (Li et al., 2023), using the YOLO11 model (Jocher and Qiu, 2024) for processing. This operation resulted in bounding boxes similar to those in object detection datasets.
@@
Next, we apply the Laplacian operation to these datasets and compute the variance of the results. Images with a variance below the threshold are discarded, as these correspond to images with a high degree of blurriness. Before cropping, we also check the size of the human bounding boxes. Images with low-resolution human bodies are rejected. After these steps, we use the bounding boxes to crop the images. When cropping, the side length of the bounding box’s longer edge is used as the side length of the cropping box, ensuring a square crop. In cases where an image contains multiple overlapping boxes, non-maximum suppression (NMS) is applied, prioritizing the box closest to the image center. The cropped images are then resized to $\times512$.
@@
Finally, we obtain approximately 440,000 cropped images, on which we measure no-reference Image Quality Assessment (IQA) metrics. The metrics used are common in image restoration, including CLIPIQA (Wang et al., 2023a), MANIQA (Yang et al., 2022), MUSIQ (Ke et al., 2021), and NIQE (Zhang et al., 2015). To balance the evaluation performance of each metric, we normalize the obtained IQA metrics using the following standardization formula:
$\text{Normalized Metrics}=\frac{1}{N}\sum_{i=1}^{N}\frac{M_{i}-\mu_{i}}{\sigma_{i}},$ (1)
@@
where $\mu_{i}$ is the mean and $\sigma_{i}$ is the standard deviation of the metrics ($M_{i}$). Since NIQE is better when its score is smaller, we first apply a negative transformation to its values. The normalized scores are then accumulated per image and sorted. The final PERSONA dataset version consists of 109,053 images with normalized metrics in the top third, and each IQA metric must exceed a predefined threshold.